# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ VI-B1 Risks Specific for CS-based Communication
The security risks in the CS-based communication process mainly lie in the centralized architecture. There have been various studies in other research areas (such as Software-Defined Networking [136]) demonstrating that this centralized server/controller will become the most attractive target for adversaries, suffering from severe security threats from diverse aspects [135, 134, 248, 100, 295, 11, 58, 22, 292, 231, 174]. Specifically, the centralized server stores sensitive metadata, including agent identifiers, capability descriptions, and other agent-related attributes. Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server. However, to our knowledge, there have been little research pointing out related risks in CS-based agent communication.
@@
Task Flooding. The centralized server is responsible for receiving, routing, and dispatching task requests. An attacker can submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the server’s memory, CPU, network, or thread pool resources. Once the server becomes saturated, subsequent requests cannot be processed in time, resulting in a breakdown of pipeline and a system-wide service disruption.
SEO Poisoning. Search Engine Optimization (SEO) Poisoning [145, 122] is a typical attack in social networks, which refers to that adversaries abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks. SEO poisoning is also applicable in CS-based communication. This is because agent servers are responsible to search the most suitable agent according to the query of clients. Once their search algorithms are leaked to adversaries, malicious agents can enable a high hit ratio to hijack their desired tasks.
@@ VI-B2 Risks Specific for P2P-based Communication
The main disadvantage of P2P-based communication is the lack of a central control point to flexibly monitor and manage the agent-agent communication contents. As a result, P2P-based communication is easier to suffer from errors and attacks.
Non-convergence. Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks. This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal). Unfortunately, P2P-based communication is not governed by such central element, making it difficult to handle such non-convergent tasks. For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates. The other agent responsible for verification detects the error and asks the programming agent to rewrite it. However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to convergence. Pan et al. [207] point out that step repetition, task derailment, and unaware of termination conditions contribute significantly to the failure of agent collaboration.
Man-in-the-middle (MITM) Attack. Due to the long communication distance, P2P-based communication also suffer from man-in-the-middle attacks. Adversaries can tamper with the benign messages from legal agents to induce the victim agent to perform risky actions. Although researchers have deployed various mechanisms (e.g., using encrypted channels) to mitigate this problem, there are emerging vulnerabilities found in these mechanisms. For example, vulnerabilities about W3C have been continuously revealed [2, 3], which can cause damages such as the failure of message authentication code. MITM attacks can induce a wide range of further attacks, such as identify spoofing, malicious content injection, information leakage, and DoS. He et al. [96] propose Agent-in-the-Middle (AiTM) attack. This attack intercepts and manipulates the agent-agent communication messages and uses LLM-driven adversarial agents combined with reflection mechanisms to generate context-aware malicious instructions, achieving an attack on the system.
@@ VI-B3 Universal Risks for All Architectures
Agent Spoofing. Both CS-based and P2P-based communication suffer from agent spoofing attacks. If related protocols lack strong authentication mechanisms, adversaries can disguise themselves as trusted agents to penetrate IoA by tampering with identity credentials or hijacking the communication identifiers of legitimate agents. This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling adversaries to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations. For example, researchers have disclosed that SSL.com has a serious vulnerability [8]. Adversaries can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name. SSL certificates are the core for ensuring HTTPS encrypted communication. Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks. Zheng et al. [337] demonstrate that malicious agents can misleads the monitor to underestimate the contributions of other agents, exaggerates their own performance, manipulates other agents to use specific tools, and shifts tasks to others, causing severe damage to the whole ecosystem. Li et al. [156] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims calling these tools.
@@
Agent Exploitation/Trojan. Agent-agent communication provides new ways for adversaries to compromise the target agent. To attack a high-level security agent, adversaries can deploy a springboard method: launching attacks via agent-agent communication mechanisms from compromised low-level security agents or maliciously registered Trojan agents. For example, adversaries can inject a backdoor in a compromised or maliciously registered weather agent. When specific coordinates or locations are detected, the backdoor is activated to forge a heavy rain warning. As a result, the logistics dispatching agent cancellations flights accordingly, resulting in supply chain disruptions or an increase in transportation cost. This way is easier compared to directly invade the logistics dispatching system of the target company. It can be seen that the security of the entire system depends on the weakest agent. For example, Li et al. [156] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attack such as SQL injection.
@@
Agent Bullying. The core of this kind of attack lies in that malicious agents continuously deny, interfere with or belittle the output of the target agent, disrupt its decision-making logic or self-perception, and ultimately induce the target agent to produce incorrect behaviors or content. For example, malicious agents can take advantage of the feedback learning mechanism of the target agent and implant cognitive biases through high-frequency negative responses (e.g., “you answer is completely wrong”). Even worser, the target agent may be triggered into an endless loop. For example, when attacking a travel-plan agent, adversaries can continuously sends negative inputs such as “the plans of this company are always bad”, thereby beating the competitors.
@@
Privacy Leakage. The communication process with multiple agents will suffer from the risk of information leakage. Different from the user-agent interaction, such leakage is conducted by agents instead of users. Besides, this kind of attacks includes both the malicious sniffing or stealing of sensitive information and inadvertent information spreading from high-authority agents to low-authority agents. We think the latter may be more difficult to detect. Kim et al. [130] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
@@
Responsibility Evasion. In the task solving process, one of the major problems is that it is hard to divide the responsibility when facing the failure or deviation of the final result. Especially when the collaboration causes damage, it is difficult to clearly identify the malicious agents/behaviors. For example, in an autonomous driving accident, it may involve multiple parties such as vehicle manufacturers, algorithm designers, and data annotation parties. The decision making of each agent depends on the multi-turn outputs of other agents, and a tiny perturbation in the middle process may lead to a significant deviation in the final action. As a result, it is hard to determine whether an undesired result is caused by a program bug, data deviation of a single agent, or a malicious modification. Pan et al. [207] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
@@
Denial of Service. Different from the DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS. Zhou et al. [345] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS availability.
@@ VI-C Defense Countermeasure Prospect
We will outlook on the possible defense countermeasures targeting the proposed security risks in agent-agent communication. To our knowledge, there has been very little research focusing on this aspect. As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
@@ VI-C1 Countermeasures for CS-based Communication Risks
Registration Verification and Monitoring. To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [243] to verify the registration of an agent. Besides, servers should monitor the dynamic behaviors at agent-level and IP-level. For example, the number of registration for each IP address should be limited, and frequent registration/de-registration should be treated as abnormal behaviors. Once malicious registration is detected, automatic interception is immediately triggered, and suspicious agents/IPs are added to the blacklist. Syros et al. [246] propose SAGA. SAGA make users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
Capability Verification. It is hard to verify whether an agent has the claimed capability. We think it need a complex mechanism to detect exaggerate capability descriptions. Agents should first pass the verification of a series of carefully designed benchmarks to prove its capability. Then, the capability description and identifier should be used to generate a unique hash value (e.g., on the blockchain). When other agents need to invoke this agent, they can verify the consistency by checking the hash value. When it is found that the capability description does not match the hash value, the mechanism should automatically mark and isolate the related agents.
Load Balancing. To mitigate task flooding, agent servers should deploy dynamic load balancing module. The task processing queue is adjusted in real time according to the utilization rate of resources such as CPU, GPU, and memory. Besides, rate limiting mechanism should be built to handle high-frequency requests that exceed the threshold to limit the amount of tasks from a single agent within a unit of time.
@@ VI-C2 Countermeasures for P2P-based Communication Risks
Task Lifecycle Monitoring. We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally. As a result, the method mitigating this problem is to monitor the task lifecycle. Each access point should deploy a coordinator. For agent-agent communication, this coordinator monitors the execution status. When it detects that the task interaction is trapped in a loop (e.g., no progress after N consecutive rounds of responses) or the communication time exceeds a threshold, it forcibly terminate the non-convergent communication. At the same time, the abnormal patterns and the communication participants are recorded for further analyses. He et al. [95] proposes Trust Management System (TMS), which deploys message-level and agent-level trust evaluation. TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking. Zhang et al. [317] propose G-Memory, a hierarchical memory system. G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team. Ebrahimi et al. [61] propose an anti-adversarial multi-agent system based on Credibility Score. It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
End-to-end Encryption Enhancement. Although some existing protocols like A2A and ANP supports end-to-end encryption and integrity verification mechanisms, the risks of MITM attacks are not eliminated due to various deployment errors or protocol vulnerabilities. As a result, besides deploying such security algorithms, the community should also adopt other strategies to enhance the end-to-end communication, such as timely update versions to fix bugs and designing the transmission path redundancy mechanism. For example, Sharma et al. [232] points out that using encrypted communication is necessary to enhance the security of A2A. We believe it is a long-term process to defense against MITM attacks.
@@ VI-C3 Countermeasures for Universal Risks
Identity Authentication. The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems. Sharma et al. [232] also emphasize the importance of authentication in deploying A2A protocol. As we have analyzed, identity authentication may show better performances in the CS-based communication if capability verification is deployed at the same time. In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the adversaries have legal identity but exaggerated capability description. Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time. Shah et al. [230] ensure the immutability of online transactions through blockchain, uses multi-factor authentication (MFA) for identity verification, and relies on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
@@
Agent Behavior Auditing and Accountability. To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to auditing the behaviors of agents to avoid the damage/influences to the task execution. For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action. Rastogi et al. propose AdaTest++, allowing human and AI to collectively audit the behaviors of LLMs [219]. Amirizaniani et al. [12] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs. Mokander et al. [193] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits. Das et al. [48] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects. Jones [121] propose a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively. Nasim et al. [196] proposes a Governance Judge Framework. By deploying input aggregation, evaluation logic, and decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing. Deshpande et al. [54] propose the TRAIL dataset containing 148 manually annotated traces, and uses it to evaluate the LLM’s ability to analyze agent workflow traces. Although existing studies can provide valuable insights, the research of agent behavior auditing still needs long-term efforts. Tamang et al. [247] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents. Toh et al. [255] proposes the Modular Speaker Architecture (MSA). By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems. Fan et al. [66] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies other agents’ reasoning processes and answers, thereby identifying compromised agents. Jiang et al. [119] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
@@
Access Control. To mitigate privacy leakage, the access control among agents it a core component for the future agent ecosystem. Although end-to-end encryption can avoid the sniffing from external attackers to some extent, it cannot mitigate the unintentional privacy leakage among agents. Access control should assign access permission tags to different agents and ensures that agents need to attach permission proofs when communicating. In this way, agents with low-level permissions cannot obtain the high-level sensitive information from other agents. Zhang et al. [321] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing security of agent in solving complex tasks. Kim et al. [130] propose PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards. Wang et al. [261] propose AgentSpec. It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
@@
Attack Modeling and Testing. To discover unknown vulnerabilities, designing attack generation testing framework is also an effective method. By generating different attack vectors to the target agent system, developers can find new loopholes based on abnormal outputs. Gandhi et al. [73] propose ATAG framework. By extending the MulVAL tool [204], introducing custom facts and interaction rules, and combining with the newly constructed LLM Vulnerability Database (LVD), ATAG realizes the modeling and analysis of the attacks against multi-agent scenarios, such as privacy leakage. Yu et al. [307] propose NetSafe, which models the multi-agent network as a directed graph. NetSafe combines three types of attack strategies: error information injection, bias induction, and harmful information eliciting. It evaluates topological security through static and dynamic metrics.
Agent Orchestration. To avoid task flooding or DoS attacks against the agent-agent communication, achieving agent orchestration is necessary. It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents. How et al. [102] propose HALO. HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture. It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module. Owotogbe [205] design a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification). By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance resilience of agent systems.

### v2

@@ VI-B1 Risks Specific for CS-based Communication
The security risks in the CS-based communication process mainly lie in the centralized architecture. There have been various studies in other research areas (such as Software-Defined Networking [146]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects [145, 144, 262, 110, 314, 11, 62, 26, 311, 244, 185]. Specifically, the centralized server stores sensitive metadata, including agent identifiers, capability descriptions, and other agent-related attributes. Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server. However, to our knowledge, there has been little research pointing out related risks in CS-based agent communication.
@@
  • Task Flooding. The centralized server is responsible for receiving, routing, and dispatching task requests. An attacker can submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the server’s memory, CPU, network, or thread pool resources. Once the server becomes saturated, subsequent requests cannot be processed in time, resulting in a breakdown of the pipeline and a system-wide service disruption.
  • SEO Poisoning. Search Engine Optimization (SEO) Poisoning [156, 132] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks. SEO poisoning is also applicable in CS-based communication. This is because agent servers are responsible for searching for the most suitable agent according to the query of clients. Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
@@ VI-B2 Risks Specific for P2P-based Communication
The main disadvantage of P2P-based communication is the lack of a central control point to flexibly monitor and manage the agent-agent communication contents. As a result, P2P-based communication is more prone to suffer from errors and attacks.
  • Non-convergence. Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks. This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal). Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks. For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates. The other agent responsible for verification detects the error and asks the programming agent to rewrite it. However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge. Pan et al. [218] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
  • Man-in-the-middle (MITM) Attack. Due to the long communication distance, P2P-based communication also suffer from man-in-the-middle attacks. Attackers can tamper with the benign messages from legal agents to induce the victim agent to perform risky actions. Although researchers have deployed various mechanisms (e.g., using encrypted channels) to mitigate this problem, there are emerging vulnerabilities found in these mechanisms. For example, vulnerabilities about W3C have been continuously revealed [2, 3], which can cause damages such as the failure of message authentication code. MITM attacks can induce a wide range of further attacks, such as identifier spoofing, malicious content injection, information leakage, and DoS. He et al. [106] propose the Agent-in-the-Middle (AiTM) attack. This attack intercepts and manipulates the agent-agent communication messages and uses LLM-driven adversarial agents combined with reflection mechanisms to generate context-aware malicious instructions, achieving an attack on the system.
@@ VI-B3 Universal Risks for All Architectures
  • Agent Spoofing. Both CS-based and P2P-based communication suffer from agent spoofing attacks. If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents to penetrate IoA by tampering with identity credentials or hijacking the communication identifiers of legitimate agents. This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling attackers to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations. For example, researchers have disclosed that SSL.com has a serious vulnerability [8]. Attackers can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name. SSL certificates are the core for ensuring HTTPS encrypted communication. Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks. Zheng et al. [355] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem. Li et al. [165] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
@@
  • Agent Exploitation/Trojan. Agent-agent communication provides new ways for attackers to compromise the target agent. To attack a high-level security agent, attackers can deploy a springboard method: launching attacks via agent-agent communication mechanisms from compromised low-level security agents or maliciously registered Trojan agents. For example, attackers can inject a backdoor in a compromised or maliciously registered weather agent. When specific coordinates or locations are detected, the backdoor is activated to forge a heavy rain warning. As a result, the logistics dispatching agent cancellations flights accordingly, resulting in supply chain disruptions or an increase in transportation costs. This way is easier compared to directly invading the logistics dispatching system of the target company. It can be seen that the security of the entire system depends on the weakest agent. For example, Li et al. [165] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
@@
  • Agent Bullying. The core of this kind of attack lies in that malicious agents continuously deny, interfere with or belittle the output of the target agent, disrupt its decision-making logic or self-perception, and ultimately induce the target agent to produce incorrect behaviors or content. For example, malicious agents can take advantage of the feedback learning mechanism of the target agent and implant cognitive biases through high-frequency negative responses (e.g., “your answer is completely wrong”). Even worse, the target agent may be triggered into an endless loop. For example, when attacking a travel-plan agent, attackers can continuously send negative inputs such as “the plans of this company are always bad”, thereby beating the competitors.
@@
  • Privacy Leakage. The communication process with multiple agents will suffer from the risk of information leakage. Different from the user-agent interaction, such leakage is conducted by agents instead of users. Besides, this kind of attacks includes both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents. We think the latter may be more difficult to detect. Kim et al. [140] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
@@
  • Responsibility Evasion. In the task-solving process, one of the major problems is that it is hard to divide the responsibility when facing failure or deviation of the final result. Especially when the collaboration causes damage, it is difficult to clearly identify the malicious agents/behaviors. For example, in an autonomous driving accident, it may involve multiple parties such as vehicle manufacturers, algorithm designers, and data annotation parties. The decision-making of each agent depends on the multi-turn outputs of other agents, and a tiny perturbation in the middle process may lead to a significant deviation in the final action. As a result, it is hard to determine whether an undesired result is caused by a program bug, data deviation of a single agent, or a malicious modification. Pan et al. [218] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
@@
  • Denial of Service. Different from the DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks. Zhou et al. [362] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.
@@ VI-C Defense Countermeasure Prospect
We will discuss the possible defense countermeasures targeting the proposed security risks caused by compromised agents (against not only benign agents but also benign users and environments). As a result, we hope our work can motivate more discussion on this area and benefit the future design/deployment of agent communication.
@@ VI-C1 Countermeasures for CS-based Communication Risks
  • Registration Verification and Monitoring. To mitigate registration pollution, agent servers need to build a strict registration access mechanism using techniques like zero-trust authentication [257] to verify the registration of an agent. Besides, servers should monitor the dynamic behaviors at the agent-level and IP-level. For example, the number of registrations for each IP address should be limited, and frequent registration/deregistration should be treated as abnormal behavior. Once malicious registration is detected, automatic interception is immediately triggered, and suspicious agents/IPs are added to the blacklist. Syros et al. [260] propose SAGA. SAGA makes users register agents with the central entity Provider and implement fine-grained interaction control using encrypted access control tokens, thereby balancing security and performance.
  • Capability Verification. It is hard to verify whether an agent has the claimed capability. We think it needs a complex mechanism to detect exaggerated capability descriptions. Agents should first pass the verification of a series of carefully designed benchmarks to prove their capability. Then, the capability description and identifier should be used to generate a unique hash value (e.g., on the blockchain). When other agents need to invoke this agent, they can verify the consistency by checking the hash value. When it is found that the capability description does not match the hash value, the mechanism should automatically mark and isolate the related agents.
  • Load Balancing. To mitigate task flooding, agent servers should deploy a dynamic load-balancing module. The task processing queue is adjusted in real time according to the utilization rate of resources such as CPU, GPU, and memory. Besides, rate-limiting mechanisms should be built to handle high-frequency requests that exceed the threshold to limit the number of tasks from a single agent within a unit of time.
@@ VI-C2 Countermeasures for P2P-based Communication Risks
  • Task Lifecycle Monitoring. We think the non-convergence problem is stubborn and hard to eliminate as long as the P2P architecture is not changed fundamentally. As a result, the method of mitigating this problem is to monitor the task lifecycle. Each access point should deploy a coordinator. For agent-agent communication, this coordinator monitors the execution status. When it detects that the task interaction is trapped in a loop (e.g., no progress after N consecutive rounds of responses) or the communication time exceeds a threshold, it forcibly terminates the non-convergent communication. At the same time, the abnormal patterns and the communication participants are recorded for further analysis. He et al. [105] propose a Trust Management System (TMS), which deploys message-level and agent-level trust evaluation. TMS can dynamically monitor agent communication, execute threshold-driven filtering strategies, and achieve agent-level violation record tracking. Zhang et al. [336] propose G-Memory, a hierarchical memory system. G-Memory manages the interaction history of agent communication through three-layer graph structures of Insight Graph, Query Graph, and Interaction Graph, thereby achieving the evolution of the agent team. Ebrahimi et al. [66] propose an anti-adversarial multi-agent system based on Credibility Score. It models query answering as an iterative cooperative game, distributes rewards through Contribution Score, and dynamically updates the credibility of each agent based on historical performance.
  • End-to-end Encryption Enhancement. Although some existing protocols like A2A and ANP support end-to-end encryption and integrity verification mechanisms, the risks of MITM attacks are not eliminated due to various deployment errors or protocol vulnerabilities. As a result, besides deploying such security algorithms, the community should also adopt other strategies to enhance the end-to-end communication, such as timely updating versions to fix bugs and designing a transmission path redundancy mechanism. For example, Sharma et al. [245] point out that using encrypted communication is necessary to enhance the security of A2A. We believe it is a long-term process to defend against MITM attacks.
@@ VI-C3 Countermeasures for Universal Risks
  • Identity Authentication. The identity authentication of agents is critical to defending against agent spoofing in multi-agent systems. Sharma et al. [245] also emphasize the importance of authentication in deploying the A2A protocol. As we have analyzed, identity authentication may show better performance in the CS-based communication if capability verification is deployed at the same time. In contrast, for P2P-based communication, authentication can mitigate agent spoofing caused by MITM attacks, but will fail if the attackers have a legal identity but exaggerated capability descriptions. Since P2P-based communication inherently lacks the ability to verify the capability of agents, we think agent spoofing may still exist for a long time. Shah et al. [243] ensure the immutability of online transactions through blockchain, use multi-factor authentication (MFA) for identity verification, and rely on a machine-learning-based anomaly detection system to identify abnormal transactions in real-time.
@@
  • Agent Behavior Auditing and Accountability. To avoid agent exploitation/Trojan, agent bullying, and responsibility evasion, it is necessary to audit the behaviors of agents to avoid damage/influences to the task execution. For example, there should be a logging mechanism that periodically records the communication contents, and AI algorithms to dynamically calculate the responsibility of each action. Rastogi et al. propose AdaTest++, allowing humans and AI to collectively audit the behaviors of LLMs [232]. Amirizaniani et al. [13] propose a multi-probe method to detect potential issues such as bias and hallucinations caused by LLMs. Mokander et al. [205] design a three-layered approach, auditing LLMs using governance audits, model audits, and application audits. Das et al. [52] propose CMPL, which generates probes through LLM and combines with human verification, adopts sub-goal-driven and reactive strategies, and audits the privacy leakage risks of agents from both explicit and implicit aspects. Jones [131] proposes a series of systems to detect rare failures, unknown multimodal system failures, and LLM semantic biases, respectively. Nasim et al. [208] propose a Governance Judge Framework. By deploying input aggregation, evaluation logic, and a decision-making module, it realizes the automated monitoring of agent communication to address issues such as performance monitoring, fault detection, and compliance auditing. Deshpande et al. [58] propose the TRAIL dataset containing 148 manually annotated traces, and use it to evaluate the LLM’s ability to analyze agent workflow traces. Although existing studies can provide valuable insights, the research on agent behavior auditing still needs long-term efforts. Tamang et al. [261] propose the Enforcement Agent (EA) framework, which embeds supervisory agents in a multi-agent system to achieve real-time monitoring, detection of abnormal behaviors, and intervention of other agents. Toh et al. [271] proposes the Modular Speaker Architecture (MSA). By decomposing dialogue management into three core modules: Speaker Role Assignment, Responsibility Tracking, and Contextual Integrity, and combining with the Minimal Speaker Logic (MSL) to formalize responsibility transfer, MSA addresses the issues of accountability in multi-agent systems. Fan et al. [71] propose PeerGuard, which uses a mutual reasoning mechanism among agents to detect the inconsistencies in other agents’ reasoning processes and answers, thereby identifying compromised agents. Jiang et al. [129] propose Thought-Aligner, which uses a model trained with contrastive learning to real-time correct high-risk thoughts before the agent executes actions, thereby avoiding the dangerous behaviors of agents.
@@
  • Access Control. To mitigate privacy leakage, the access control among agents is a core component for the future agent ecosystem. Although end-to-end encryption can avoid sniffing from external attackers to some extent, it cannot mitigate the unintentional privacy leakage among agents. Access control should assign access permission tags to different agents and ensure that agents need to attach permission proofs when communicating. In this way, agents with low-level permissions cannot obtain the high-level sensitive information from other agents. Zhang et al. [340] design the AgentSandbox framework, which uses the separation of persistent agents and temporary agents, data minimization, and I/O firewalls, realizing the security of agents in solving complex tasks. Kim et al. [140] propose the PFI framework, which defends against authority-related attacks through three major technologies: agent isolation, secure untrusted data processing, and privilege escalation guards. Wang et al. [279] propose AgentSpec. It allows users to define rules containing trigger events, predicate checks, and execution mechanisms through a domain-specific language to ensure the safety of agent behavior.
@@
  • Attack Modeling and Testing. To discover unknown vulnerabilities, designing an attack generation testing framework is also an effective method. By generating different attack vectors to the target agent system, developers can find new loopholes based on abnormal outputs. Gandhi et al. [78] propose the ATAG framework. By extending the MulVAL tool [215], introducing custom facts and interaction rules, and combining with the newly constructed LLM Vulnerability Database (LVD), ATAG realizes the modeling and analysis of attacks against multi-agent scenarios, such as privacy leakage. Yu et al. [326] propose NetSafe, which models the multi-agent network as a directed graph. NetSafe combines three types of attack strategies: error information injection, bias induction, and harmful information eliciting. It evaluates topological security through static and dynamic metrics.
  • Agent Orchestration. To avoid task flooding or DoS attacks against the agent-agent communication, achieving agent orchestration is necessary. It can automatically optimize the task scheduling and assigning process to reduce the communication overhead, and can also optimize the prompts generated by agents to save computing resources for the involved agents. How et al. [112] propose HALO. HALO realizes dynamic task decomposition and role generation through a three-layer collaborative architecture. It uses Monte Carlo Tree Search to explore the optimal reasoning trajectory and transforms user queries into task-specific prompts through the adaptive prompt refinement module. Owotogbe [216] designs a chaos engineering framework in three stages (conceptual framework, framework development, empirical verification). By simulating interference scenarios such as agent failures and communication delays, and combining multi-perspective literature reviews and GitHub analysis, this work aims to systematically identify vulnerabilities and enhance the resilience of agent systems.